OpenStudy (anonymous):

Use the method of substitution to solve the system of linear equations. 7x+4y=7/6 y=6x-1

OpenStudy (anonymous):

just substitute y=6x-1 in the first equation 7x+4 (6x-1) = 7/6 solve for x then substitute for x in y=6x-1

OpenStudy (anonymous):

I can't solve for x, thats the problem.

OpenStudy (anonymous):

\[7x + 4(6x-1)= \frac{ 7 }{ 6}\] \[7x+24x-4=\frac{ 7 }{ 6 }\] \[31x = \frac{ 7 }{ 6 } + 4\] \[31x = \frac{ 31 }{ 6 } * both sides \frac{ 1 }{ 31 }\] \[x = \frac{ 1 }{ 6 }\] \[y = 6 (\frac{ 1 }{ 6 }) -1 = 0\]

OpenStudy (anonymous):

0,0?

OpenStudy (anonymous):

1/6 , 0